### v1

# A First Look at Bugs in LLM Inference Engines

Abstract.
Large language model-specific inference engines (in short as LLM inference engines) have become a fundamental component of modern AI infrastructure, enabling the deployment of LLM-powered applications (LLM apps) across cloud and local devices. Despite their critical role, LLM inference engines are prone to bugs due to the immense resource demands of LLMs and the complexities of cross-platform compatibility. However, a systematic understanding of these bugs remains lacking. To bridge this gap, we present the first empirical study on bugs in LLM inference engines. We mine official repositories of 5 widely adopted LLM inference engines, constructing a comprehensive dataset of 929 real-world bugs. Through a rigorous open coding process, we analyze these bugs to uncover their symptoms, root causes, and commonality. Our findings reveal six major bug symptoms and a taxonomy of 28 root causes, shedding light on the key challenges in bug detection and location within LLM inference engines. Based on these insights, we propose a series of actionable implications for researchers, inference engine vendors, and LLM app developers.
copyright: nonejournalyear: 2025
1. Introduction
Large language model-specific inference engine (abbreviated as LLM inference engine) has emerged as a vital component of modern artificial intelligence infrastructure. These engines are specially optimized for LLM inference, offering LLM-specific optimizations tailored to reduce resource consumption and enhance inference speed. LLM inference engines have significantly boosted the adoption of LLM-powered applications (LLM apps). While LLM inference engines deliver fast, high-quality responses in cloud environments (cha, 2024), they also support local deployment on personal devices like Windows, Linux, and MacOS PCs for personal LLM apps such as Ollama (oll, 2024).
LLM inference engine is complexly constructed and highly bug-prone due to the immense resource demands of LLMs and the challenges of cross-platform compatibility. First, due to the vast number of parameters in LLMs, LLM inference is both compute-intensive and memory-hungry. To enhance inference efficiency, LLM inference engines employ numerous complex optimizations to reduce the memory consumption and enhance the inference latency, such as paged attention, continuous batching, and speculative decoding. Second, the engines should adapt to a wide range of application settings. They need to support various LLM architectures across different environments, including multi-node clusters, single nodes with multi-GPUs, single GPU servers, PCs, mobile devices, IoT devices, and browsers. The combination of the complex optimizations and diverse environment adoption potentially leads to more bugs, which can severely impact the availability, performance, and user experience of LLM apps.
Bugs in AI frameworks have been extensively studied. Previous work has empirically investigated bugs of deep learning (DL) frameworks across stages, including the learning stage (Islam et al., 2019; Zhang et al., 2018, 2020; Humbatova et al., 2020), the compiling stage (Shen et al., 2021; Guan et al., 2023), and the deployment stage (Chen et al., 2020, 2021). However, a systematic understanding and characterization of bugs in LLM inference engines remains lacking. This knowledge is urgently needed as LLMs have become the dominant AI paradigm.
To bridge this gap, we present the first empirical study on bugs in LLM inference engines. Given the increasing prevalence of LLM apps, this research is of great significance. Specifically, we aim to address three research questions (RQs):
(RQ1) Symptoms: What are the symptoms of bugs and their distribution? Understanding the symptoms helps clarify the impacts of bugs and guides the efficient setting of test oracles.
(RQ2) Root Causes: What are the root causes of bugs and their distribution? The root causes help identify vulnerabilities and guide bug detection and localization.
(RQ3) Commonality: Do bugs follow a common pattern across engines? Investigating commonality reveals cross-engine generalization of bug patterns and shared vulnerabilities to guide the testing methods.
To answer the RQs, we carefully select five representative LLM inference engines for our analysis, including Llama.cpp (lla, 2024b), vLLM (vll, 2025), DeepSpeed (dee, 2025), Mlc-llm (team, 2023), and TensorRT-llm (ten, 2025). These engines are commonly used in real-world LLM apps and exhibit diversity in their implementations, ensuring the validity and generalizability of our results. We gather all the bug issues from the official repositories of these engines as of December 1, 2024. Following the open coding procedure (Seaman, 1999), we review error reports, source code, developer discussions, and related pull requests (PRs) of issues to uncover the symptoms, root causes, and distribution of bugs, constructing a dataset with 929 manually labeled real-world bugs. Based on the dataset, we identify six symptom types and construct a taxonomy of 28 specific root causes. Our main findings include that (1) Under undetermined LLM output, five factors (i.e., mis-encoded characters, incoherent semantics, inconsistent outputs, abnormal lengths, and repeated substrings) provide insights for oracle setting; (2) Testing LLM inference engines requires diverse oracles beyond mere crash detection, as over 35% of bugs manifest as non-crash symptoms. (3) We identify actionable diagnostic patterns for the observed bug symptoms. For instance, crashes, feature failures, and system hangs necessitate checks on functionality, environment, and configuration. (4) Bug symptoms and root causes show a strong correlation across engines, with correlation coefficient greater than 0.5, indicating common bug patterns despite different engine implementations.
In summary, we make the following main contributions:
  • To the best of our knowledge, we present the first empirical study on bugs in LLM inference engines.
  • We present taxonomies of bug symptoms and root causes in LLM inference engines, which provide practical insights into bug detection and location.
  • We make our carefully collected dataset from our work public available for further research [^1] https://github.com/infbug/bugs-in-LLM-inference-engines.
2. Background
In this section, we introduce the deployment workflow of LLMs and the main components of LLM inference engines.
2.1. LLM Deployment
Figure 1. Workflow of LLM deployment.
LLM apps have gained widespread use, such as cloud-based chatbots (cha, 2024), and personal AI assistants deployed on user devices (app, 2024; goo, 2024). In these apps, LLM inference engines are crucial components for enhancing inference efficiency and adapting LLMs to various environments. Generally, the workflow of LLM deployment falls into three phases, including engine setup, model conversion, and inference/serving, as shown in Fig. 1.
Engine setup aims to deploy LLM inference engines onto target devices (e.g., distributed servers, PCs, and phones), either through source code compilation or pre-compiled binaries. This phase involves installing necessary software modules and ensuring hardware compatibility with specific setup configurations.
Model conversion adapts models for specific engines, platforms, and devices through steps like format conversion, parameter optimization, and packaging. Pre-trained LLMs are adjusted for specific conversion configurations, and efficiency optimizations like quantization and sharding are applied based on the underlying environments. Given the high resource and performance demands of LLM inference, this process is crucial for ensuring the quality of LLM inference services.
Inference/serving aims to initialize services, load models, and process inference requests correctly and efficiently on target devices. Developers configure the LLM as a service, allowing users to set job-specific configurations (such as the model, backend, and inference strategies), submit prompts, and receive LLM responses.
2.2. LLM Inference Engine
Figure 2. Architecture of LLM inference engine
The primary design goals of LLM inference engines are compatibility and efficiency. First, compatibility ensures that various series of LLM run smoothly across diverse environments. Engines like vLLM (vll, 2025) and DeepSpeed (dee, 2025) are tailored for diverse cloud-based distributed inference scenarios, while engines like Llama.cpp (lla, 2024b) and Mlc-llm (team, 2023) focus on compatibility with heterogeneous personal devices and backends. Second, efficiency optimization focuses on integrating algorithms that maximize resource utilization and minimize inference latency under varying resource conditions. Engines generally achieve efficiency by optimizing model conversion, inference strategies, operator implementation, and resource management.
As shown in Fig. 2, LLM inference engines typically support dozens of popular LLM series, such as Llama (lla, 2024a) and Gemma (gem, 2024). Some of the engines (lla, 2024b; team, 2023; ten, 2025) provide generally supported model formats and conversion tools for user-customized LLMs. They accommodate a wide range of devices, including distributed servers, single-GPU servers, PCs, smartphones, and IoT devices. Some of them (lla, 2024b; team, 2023) are also compatible with various platforms like Linux, Windows, and Android, as well as backends such as CUDA, Vulkan, and Metal. Next, we discuss the major components of the engines.
These engines often provide setup tools for custom engine construction and deployment in the engine setup phase and offer model conversion tools for the model conversion phase. In the inference/serving phase, users can interact with LLMs via command lines or inference services. Upon receiving a start command, the engine initializes the LLM inference service based on the configurations and loads the necessary LLMs onto the selected backend. When the service receives a user request with specific prompts and job configurations, the engine allocates resources, creates a job, checks the chosen model, and performs auto-regressive token generation to respond to users. This process involves several key components, as shown in Fig. 2.
The engine initializer processes service configurations to set up the inference environment, including loading dependent modules, configuring parameters, coordinating with the model loader to prepare LLMs, and opening the service port to await inference requests. The model loader is responsible for loading the specified converted LLM onto the chosen backend.
For each inference request, the request parser processes the input prompts and job configurations, formatting them for the job manager. The job manager handles concurrent jobs and their contexts. For each parsed request, the job manager creates a job from the configurations and collaborates with the model loader and resource manager to ensure smooth inference. For each job, the input processor tokenizes the prompts and applies LLM-specific features, such as grammar constraints to enforce specific formats. The engine then generates tokens auto-regressively based on backend-specific operators. Finally, the output processor detokenizes the output tokens, converts them to the specified output format, and sends the response to the user.
The resource manager employs optimized algorithms to manage critical computational and memory resources for LLM inference, such as the key-value cache (KV cache), which stores and reuses the generated key and value vectors to reduce redundant calculations. The backend adapter manages API interactions with the hardware, supporting the resource manager’s operations. The communicator is tailored for distributed environments, facilitating message passing and status synchronization between servers. Engines also typically include monitors and loggers to track metrics and logs during inference, as well as testing and benchmarking tools to evaluate inference accuracy and performance.
3. Methodology
Fig. 3 shows an overview of our methodology. We begin by selecting representative LLM inference engines, then collect relevant bugs, and finally perform manual labeling to characterize them.
Figure 3. Overview of methodology.
3.1. Engine Selection
With the rising popularity of LLMs, numerous LLM inference engines have been developed, making it impractical to consider them all. Therefore, we select representative ones for our study. Notably, we focus on LLM-specific inference engines, excluding those designed for general DL model inference.
We begin with a widely recognized leaderboard of state-of-the-art LLM training and inference projects actively maintained by the community (awe, 2025). Since our study focuses on inference-related bugs, we first exclude engines dedicated to LLM training. We also remove closed-source engines due to their lack of accessible bug reports. From the remaining candidates, we select the top five LLM inference engines based on the number of stars in their official GitHub repositories, a widely used indicator of popularity (Liu et al., 2024; Li et al., 2020). As a result, we adopt Llama.cpp (lla, 2024b), vLLM (vll, 2025), DeepSpeed (dee, 2025), MLC-LLM (team, 2023), and TensorRT-LLM (ten, 2025). The details of these engines are summarized in Table 1.
Table 1. Details of selected LLM inference engines.
Engine Provider Targeted Env. GitHub Stars
Llama.cpp Ggml Universal 76,157
vLLM UCB Cloud 40,876
DeepSpeed Microsoft Cloud 37,282
Mlc-llm CMU Universal 20,150
TensorRT-llm Nvidia Cloud/Edge 9,641
These engines are representative for several reasons. First, they have received between 9,641 and 76,157 GitHub stars, indicating strong popularity. Second, they have been adopted by major organizations. For example, vLLM serves as the backend for Hugging Face, DeepSpeed is integrated into Microsoft’s infrastructure, and TensorRT-LLM is utilized by Nvidia. Third, they show diversity in targeted environments. Llama.cpp and MLC-LLM are designed for universal deployment, supporting various operating systems (such as Linux, Windows, macOS) and platforms, including PCs, mobile devices, and IoT devices. vLLM and DeepSpeed are optimized for cloud environments, while TensorRT-LLM supports both cloud-based deployments and edge servers. Finally, these engines incorporate cutting-edge features in LLM inference, such as continuous batching, paged attention, speculative decoding, and machine learning compilation.
3.2. Bug Collection
To collect bugs from the selected LLM inference engines, we focus on their official GitHub repositories. Specifically, we first follow previous work (Chen et al., 2023; Zhang et al., 2018) to leverage the GitHub issue tracking system to gather potential bug reports for each engine. We then refine the collected data and exclude invalid reports.
3.2.1. Identifying Potential Bug Reports.
Since not all issues are related to bugs, we follow previous work (Chen et al., 2021) by using repository-specific labels to identify potential bug reports. We thoroughly review each repository to understand its labeling system. For Llama.cpp, vLLM, Mlc-llm, and TensorRT-llm, ‘label:bug’ indicates bug reports. For DeepSpeed, which supports both training and inference, bug reports related to inference are identified only when both ‘label:bug’ and ‘label:inference’ are used together. We then filter the issues based on these labels for each engine. Additionally, following prior studies (Chen et al., 2021), we exclude open issues, as they are typically unresolved and may lack clear insights into the root causes. After this process, we obtain 184, 1200, 187, 551, and 366 potential bug reports for Llama.cpp, vLLM, DeepSpeed, Mlc-llm, and TensorRT-llm, respectively, as of December 1, 2024.
3.2.2. Refining Dataset
We observe a significant number of false positives in the collected bug reports, as the bug-related tags can be misapplied. To address this, manual filtering is necessary to remove these false positives. However, manually analyzing all the issues is a challenging task. Therefore, we adopt a sampling approach following previous work (Zhang et al., 2019; Chen et al., 2020), randomly selecting a subset of the bug reports while ensuring a 95% confidence level and a 5% confidence interval. This sampling process is applied separately to each engine, resulting in 125, 292, 126, 227, and 188 bug reports for Llama.cpp, vLLM, DeepSpeed, Mlc-llm, and TensorRT-llm, respectively.
After sampling, we manually refine the dataset by excluding reports that meet any of the following criteria:
  • Not a bug. Issues that only contain feature requests or general inquiries are excluded, as they do not represent bugs.
  • Not a bug of the engine itself. Issues arising from user misunderstandings or misuse of the engines are excluded, as they do not reflect inherent engine bugs.
  • No informative comments. Issues without substantial or informative comments are excluded, as they do not provide sufficient details for analysis.
After this manual refinement, we are left with a dataset of 929 real-world bugs related to LLM inference engines. The final counts are 124, 290, 121, 214, and 180 reports for Llama.cpp, vLLM, DeepSpeed, Mlc-llm, and TensorRT-llm, respectively.
3.3. Manual Labeling
We manually analyze the 929 bugs to identify their symptoms and root causes, constructing corresponding taxonomies for LLM inference engine bugs. Following previous work (Chen et al., 2021; Guan et al., 2023), the manual analysis involves two key steps: initially constructing the taxonomies through a pilot labeling process, followed by a reliability analysis to evaluate the taxonomies. The entire manual analysis takes over 500 man-hours.
3.3.1. Pilot Labeling
We randomly select 50% of the dataset, totaling 465 issues, for pilot labeling. As the first study to characterize bugs in LLM inference engines, we develop the taxonomies for bug symptoms and root causes from scratch, following the open coding procedure (Seaman, 1999).
First, two authors, each with over two years of experience in LLM-based application development, act as inspectors and jointly develop pilot taxonomies for bug symptoms and root causes. To fully understand the content of each issue, they review the title, description, code snippets, comments, answers, labels, associated PRs, and any referenced URLs. They then assign brief phrases as initial codes to capture the key features of each issue.
For symptoms, the inspectors focus on what the bug reporter observes and at which stage of LLM deployment the bug occurs. This information is often found in the issue descriptions and call stacks provided by reporters. For example, a reporter states that ‘loading a gptbigcode model that has been quantized with gptq fails’ in issue #8116 of vLLM. The inspectors classify the symptom as a crash occurring in the inference/serving stage.
In contrast, identifying root causes is more challenging and often requires in-depth analysis. Inspectors determine which part of the code failed and how by examining call stacks, bug reproduction experiments, confirmed comments from developers, issue descriptions, and associated PRs. If the relevant code is not within a modular component, they infer its component based on its functionality. For instance, in issue #3821 of vLLM, the associated PR #4021 states ‘avoid allocating a CUDA context for every GPU’, and modifies the allocation of the cache dictionary in its commits. Based on this, inspectors conclude that the bug originated from incorrect allocation in the engine’s resource manager.
After generating the initial codes for the issues, inspectors group similar codes into categories. The grouping process is iterative, with inspectors continuously reviewing the categories and issues to refine the taxonomies. Bugs associated with multiple symptoms or root causes are assigned to all relevant categories. In case of disagreements, another author, who has over two years of LLM inference engine deployment experience, acts as an arbitrator and joins the discussion to help reach consensus. The arbitrator, as well as two inspectors, finally approve all categories in the symptom taxonomy and the root cause taxonomy.
3.3.2. Reliability Analysis
This step aims to validate the reliability of the pilot taxonomies using the remaining 464 issues.
The two inspectors independently classify each issue according to the symptom and root cause taxonomies, following the coding scheme established during pilot labeling. If an inspector is unsure how to classify an issue, it is marked as ‘Pending.’
To evaluate the reliability of our coding schema and procedure, we measure the inter-rater agreement between the two inspectors using Cohen’s Kappa ($\kappa$(Cohen, 1960), a widely adopted metric in empirical software engineering studies (Chen et al., 2021, 2023). The computed $\kappa$ values for symptom and root cause labeling are 0.927 and 0.873, respectively, indicating almost perfect agreement and confirming the reliability of our coding schema.
The arbitrator participates in discussions to resolve any labeling conflicts between the inspectors. For issues marked as ‘Pending,’ they collectively determine whether new categories are needed. Ultimately, no new category of symptoms and root causes is introduced. All pending issues, comprising only 0.9% of the total analyzed cases, are successfully classified within the existing taxonomies, demonstrating the saturation of the taxonomies. The final labeling results are reviewed and approved by all participants.
4. RQ1: Symptoms
To clarify the impact of bugs and provide insights for test oracle, we explore the bug symptoms and their distribution of LLM inference engines.
4.1. Symptom Classification
As shown in Table 2, we classify the observed symptoms into six unique categories, including Crash (S1), Unexpected Output (S2), Feature Failure (S3), Abnormal Performance (S4), System Hang (S5), and Silent Error (S6). These symptoms vary in distribution across three phases of LLM deployment workflow (i.e., engine setup, model conversion, and inference/serving), as shown in Table 3. Next, we illustrate the symptoms by phase.
{mdframed}
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Finding 1: Six bug symptoms are identified, including crash, unexpected output, feature failure, abnormal performance, system hang, and silent error.
4.1.1. Engine Setup
During engine setup, we identify 136 bugs (15% of total), covering three symptoms including S1, S5, and S2.
Specifically, most (97%) symptoms in this phase are Crash (S1) with 132 instances, indicating abrupt termination during engine source code compilation or pre-compiled binary execution. For example, in Llama.cpp issue #279, the engine compilation process terminated unexpectedly due to missing Apple Accelerate header files. Among the remaining four bugs, three exhibit System Hang (S5), where the process becomes unresponsive without crashing. For example, in vLLM issue #3601, the construction of the docker image was stuck for a very long time. The last remaining bug falls into Unexpected Output (S2), where an incorrect warning occurred during the engine construction in DeepSpeed issue #6486.
{mdframed}
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Finding 2: Most (97%) bugs in engine setup phase lead to a crash.
Table 2. Symptom Categories in LLM Inference Engines.
ID Symptom Description
S1 Crash Engines terminate unexpectedly without recovery capability.
S2 Unexpected Output Engines produce unexpected or incorrect outputs.
S3 Feature Failure Engines remain operational but fail to perform specific features.
S4 Abnormal Performance Engines significantly deviate from expected performance (e.g., resource usage and response latency).
S5 System Hang Engines remain operational but become unresponsive.
S6 Silent Error Engines experience internal anomalies without error detection.
Table 3. Number of bug symptoms in different phases.
S1 S2 S3 S4 S5 S6 Total
Engine Setup 132 1 0 0 3 0 136
Model Conversion 108 0 14 0 2 8 132
Inference/Serving 363 121 93 53 30 4 664
Total 603 122 107 53 35 12 932
4.1.2. Model Conversion
We identify 132 bugs (14% of total) in the model conversion phase, covering four symptoms including S1, S5, S3, and S6.
Specifically, the majority (82%) symptoms are classified into Crash (S1) with 108 instances, where the model conversion process terminates abruptly. For instance, in TensorRT-llm Issue #899, the conversion of LLM to ‘SmoothQuant’ format failed due to mismatched shape of tensors. Two bugs show System Hang (S5), where the model quantization process stalls without termination (e.g. TensorRT-llm issue #1356).
22 bugs result in the generation of incorrect model files. Among these bugs, 14 are classified as Feature Failure (S3), where the generated complete model files are incomplete or misconfigured under specific settings. For example, in TensorRT-LLM issue #1245, the conversion tools failed to generate complete ‘INT4’ model file under the AWQ quantization settings. The other 8 bugs are identified as Silent Error (S6), where conversion tools generate seemingly complete yet incorrect model files without any warnings. For example, in Llama.cpp issue #9245, although the input model file was broken, the engine generated model files without warnings. This highlights the need for thorough checks of generated model files to prevent incorrect models from causing severe inference errors.
{mdframed}
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Finding 3: In model conversion phase, 83% of bugs manifest as failures to generate model files (S1, S5) while 17% produce incorrect files (S3, S6).
4.1.3. Inference/Serving
We observe 664 bugs (71% of total) in the inference/serving phase, covering all six identified symptoms.
Most (59%) bugs cause the engine to stop functioning, involving symptoms of S1 and S5. Crash (S1) is the most common symptom, with 363 instances involving initialization failures or irrecoverable service termination. For example, in vLLM issue #8952, the server crashed when receiving requests with configurations like ‘response_format’ or ‘guided_json’ System Hang (S5) occurs in 30 cases, where the service becomes unresponsive without crashing, as seen in Mlc-llm issue #306, where API calls hang during prefill function execution.
93 bugs make engines exhibit functional anomalies while remaining operational, which are classified into Feature Failure (S3). Common issues include abnormal API response formats or task execution failures with user-specific configurations. For instance, in vLLM’s issue #8947, the engine mishandles the configuration priority logic.
Bugs can also happen even if engines generate output tokens, exhibiting as S2, S4, and S5. Unexpected Output (S2), with 121 instances, occurs when the service produces user-unexpected tokens or metric values (e.g., probability distribution). For example, in the DeepSpeed issue #3285, the tensor computation results of the engine differ from those of the original PyTorch model. showing parameter size consistency with PyTorch but logits deviations Abnormal Performance (S4), with 53 bugs, manifests as either excessively high or low performance metrics (e.g., latency, throughput, and resource utilization). For instance, issue #2488 of DeepSpeed reports an unexpected increased text generation latency. Silent Error (S6), with 4 bugs, involves hidden code defects that may disrupt computations. For example, the issue #6030 of vLLM reports that an if-else condition branches contained identical code blocks.
{mdframed}
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Finding 4: The inference/serving phase has the highest bug proportion (71%) and exhibits all six symptom types.
Figure 4. Distribution of unexpected factors.
Note that a key challenge in testing LLM inference is setting test oracles for their inherently undetermined output tokens. Therefore, we further analyze bugs related to unexpected output tokens to determine what constitutes ‘unexpected’ outputs.
Following the labeling process in section 3, among the 121 ‘Unexpected Output’ bugs in inference/serving phase, we further identify 102 bugs [^2] The other 19 bugs are not associated with output tokens, such as DeepSpeed issue #3285. with unexpected output tokens, covering 5 factors, whose distribution is shown in Fig. 4. The most (36%) observed factor is unexpected Character (UT1), with 37 bugs, which indicates that the output tokens of LLMs involve massive garbled text (e.g. Llama.cpp issue #1735). Semantic Coherence (UT2) is the second most (20%) used factor, found in 20 cases, implying that the outputs of LLM are meaningless (e.g. Llama.cpp issue #3782). Inconsistent Output (UT3) indicates that even after disabling all randomness settings, the LLM outputs still vary across different optimization algorithms or runtime environments. This factor accounts for 17%, with 17 instances. For instance, in vLLM issue #590, output tokens differ between the original model and the model served using vLLM. 15% of bugs, found in 15 issues, are identified by unexpected Length (UT4) of output tokens, occur when outputs are unusually short or excessively long, never stopping, exemplified by vLLM issue #9448. Repetition (UT5) issues, seen in 13 cases, involve excessive duplication of strings. For example, in DeepSpeed issue #4903, the generated tokens abnormally repeat the ‘e’ character.
{mdframed}
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Finding 5: Despite non-deterministic outputs of LLMs, five factors (i.e. character, semantic coherence, inconsistent output, length, and repetition) provide insights for oracle settings.
4.2. Symptom Distribution
Overall, Crash (S1) is the most prevalent (65%) symptom across all three phases. Unexpected Output (S2) ranks second at 13%, primarily occurring during the inference/serving phase. Feature Failure (S3) follows at 11%, with Performance Anomalies at 6%, System Hangs at 4%, and Silent Errors at 1%. Among these symptoms, crashes are commonly used for bug detection in DL libraries (Deng et al., 2023, 2024) due to their distinct characteristics and easy detection through core dumps or stack traces. Moreover, unexpected outputs are the second most common symptoms, comprising nearly 20% during the inference/serving phase. This underscores the need for suitable test oracles for LLM outputs and highlights the significance of Finding 4.
{mdframed}
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Finding 6: While crashes are the most common, over 35% of bugs show other symptoms, necessitating diverse oracles beyond crashes for LLM inference engine testing.
5. RQ2: Root Causes
To identify the vulnerabilities and obtain insights into bug detection and localization, we study the root causes and their corresponding components of LLM inference engines.
5.1. Root Cause Classification
Following the labeling process in section 3, we identify 28 specific leaf categories of root causes (e.g., Incompatible Model) and group them into 5 inner categories linked to key factors in LLM inference process (including In-/Output, Configuration, Algorithm, Environment, and Resource). Fig. 5 shows our hierarchical taxonomy. For each category, the top-left circle number indicates the bug count it causes. We discuss and exemplify each inner category below.
{mdframed}
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Finding 7: We develop a taxonomy of 28 specific root causes of bugs in LLM inference engines, indicating the diversity of bug sources in these complex engines.
Figure 5. Taxonomy of root causes of bugs in LLM inference engines. The top-left circle numbers indicate bug counts.
5.1.1. [RA] In-/Output
In-/output processing is crucial for user interaction during LLM inference. The engine receives and parses user input (including job configurations and prompts), decodes characters, and performs tokenization to prepare for inference. After inference, it decodes the generated tokens, formats the output, and returns it. Engines support various user interaction methods, such as HTTP requests and the command line. We observe 41 cases originating from in-/output processing, accounting for 4% of cases, spanning six leaf categories. Next, we detail each root cause following the processing flow.
For user input, inference engines must parse the formatted job configurations and prompts. The Incorrect Request Parsing (RA.2), with 10 cases, indicates a failure to validate and parse requests, or provide feedback for malformed inputs. For example, in Llama.cpp issue #3766, a request with malformed ‘system_prompt’ configuration caused a system hang without warning due to inadequate request validation. Additionally, in this process, engines must avoid Incorrect Character Encoding (RA.6), with 2 cases, to accurately encode the characters of prompts. RA.6 highlights the mishandling of incompatibilities between the engine’s and users’ character encodings. For instance, in Mlc-llm issue #804, the engine struggled with parsing some Unicode/UTF-8 characters of user prompts correctly.
After parsing requests, engines will tokenize the prompt for subsequent inference. Incorrect Tokenization (RA.1), with 12 cases, indicates the use of misconfigured tokenizer or incorrect tokenization algorithm, leading to wrong token splits. For instance, in Llama.cpp issue #290, the tokenizer improperly tokenized strings like ‘china’. RA.1 is the primary cause of in-/output relevant bugs, underscoring the need for meticulous attention to tokenizer checks, handling, and error tolerance. Moreover, engines face challenges in parsing multimodal data, often encountering issues with format parsing and dimension alignment for multimodal inputs, i.e., Insufficient Multimodal Support (RA.3) with 9 cases. For example, in vLLM issue #9128, the engine failed to correctly merge image and text embeddings.
After token generation, abnormal outputs can occur if the output pipeline misconfigures detokenization settings or algorithms, i.e., Incorrect Detokenization (RA.5) with 3 cases. For example, in Llama.cpp issue #4036, the output Unicode characters displayed correctly in the command line but failed in server mode. Incorrect Formatting (RA.4) of user-specified output, with 5 cases, can also lead to bugs, as demonstrated in vLLM issue #5334, where the requested log probability of tokens was not returned.
{mdframed}
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Finding 8: Although in-/output relevant bugs are less frequent (comprising 4% of bugs), they significantly impact user interaction, necessitating careful checking, processing, and error handling of character encoding, tokenizer, and user-specific formatting.
5.1.2. [RB] Configuration
Configurations are crucial throughout the entire LLM deployment workflow. Engine setup requires configurations such as hardware acceleration support. Then, model conversion requires settings such as precision. Next, inference/serving involves service configurations to initiate LLM services, specifying factors like the path of LLMs. After that, users can also specify job configurations, such as the maximum context length. We identify 242 cases (23% of total) caused by configuration issues, divided into six categories. This vast number of bugs highlights the challenges in configuration setup and management in LLM inference. Next, we discuss the detailed root causes linked to configurations.
Misconfiguration (RB.1) indicates the incorrect setup or implementation of individual configurations or their combinations. For instance, in DeepSpeed issue #3628, the engine failed to split the LLM under configurations of tensor parallelism. This category, with 105 cases, is highly bug-prone due to the engine’s extensive parameter support. This underscores the necessity of rigorous testing to ensure each configuration’s correctness and compatibility with others.
After correct configuration setup, 82 cases are linked to the configured LLMs, including (RB.2) and (RB.4). 65 cases indicate incorrect configuration and loading of LLMs, referred to as Incompatible Model (RB.2). For example, in TensorRT-llm issue #1634, quantization of Llama-3-70B-Instruct with w4a16 parameters failed because the engine lacked support for the Llama-3 architecture. Engines require individual adaptations for each LLM, which can lead to potential oversights for new models. This highlights the necessity for comprehensive engine compatibility testing for new LLMs. Additionally, 17 cases exhibit incorrect precision settings impacting model conversion, loading, and inference, referred to as Mismatched Precision (RB.4). For instance, in TensorRT-llm issue #1303, using INT8 precision during model conversion alongside a BF16 precision engine resulted in incorrect output tokens.
Moreover, to improve resource utilization and accelerate inference, engines are optimized for handling multiple requests under diverse concurrency configurations, which is also the major source of bugs, with 42 cases. Concurrency Issues (RB.5), with 39 cases, implies the incorrect management of concurrent workers. For instance, in Llama.cpp issue #3537, the static variable ‘mu’ used by the ‘Mirostat’ sampler was shared across different workers, leading to state conflicts. This implies the need for engines to carefully address potential concurrency issues, including shared variables, atomic operations, and resource contention. Moreover, Incorrect Job Context Management (RB.6), with 3 cases, refers to errors in handling context and configurations for multiple requests. For instance, in TensorRT-LLM issue #1447, the engine fails to process requests with and without the ‘SamplingConfig’ parameter simultaneously.
Moreover, Document Issues (RB.5), with 13 cases, refer to inaccurate or outdated descriptions of usage and configurations in documentation. For example, DeepSpeed issue #4039 highlights the absence of inference configuration details for MoE models.
{mdframed}
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Finding 9: Managing configurations is challenging for LLM inference engines, accounting for 23% of bugs across six root causes. Engines should prioritize comprehensive compatibility testing and timely updates for configuration combinations, model adaptation, mixed precision, concurrency conditions, multitasking, and documentation.
5.1.3. [RC] Functionality
Correct functionality implementation forms the operational backbone of LLM inference engines. We identify 432 functionality-related cases (41 % of total), spanning nine categories. This substantial number highlights the challenges in balancing mathematical accuracy, framework compatibility, and performance optimization. Next, we explain each root cause through examples.
Bug can originate from root causes that are common in DL libraries. Incorrect Algorithm Implementation (RC.1) refers to problematic functional implementation, such as the optimization strategies and computing kernels. For instance, in vLLM issue #590, an erroneous implementation of ‘RoPE’ algorithm within attention operators caused unexpected output tokens. This represents the most (43%) prevalent functionality issue with 184 cases, emphasizing the complexity of functionalities in LLM inference engines. Misused API (RC.2), with 75 cases, refers to incorrect usage of either internal or external APIs. This includes (1) failing to use an API or using it redundantly, and (2) using an incorrect API name or arguments. For example, in vLLM issue #9059, directly accessing a non-existent ’vocab’ attribute in ’MistralTokenizer’ instead of using ’get_vocab()’ caused a bug. Mismatched Shape (RC.3), with 53 cases, occurs when tensor dimensions violate operator requirements. For instance, in TensorRT-LLM issue #2150, a ‘5D’ tensor input to conv2d layers (which expect ‘3D/4D’ inputs) led to a crash.
Bugs can stem from root causes that are widespread in software. Incorrect Exception Handling (RC.4), with 37 cases, denotes failures to properly handle the exceptions. For example, in Llama.cpp issue #1768, the engine failed to raise exceptions for out-of-bounds input, resulting in a crash. Incorrect Conditional Logic (RC.5), with 24 cases, involves incorrect handling of condition branches. For instance, in Mlc-llm issue #2361, the engine failed to handle the condition where ‘committed_tokens’ is empty during the dialog initialization phase, leading to a crash. Incorrect Type (RC.6), with 24 cases, stems from type system violations. For example, Deepspeed issue #2183 mistakenly passed ‘ABCMeta’ class objects instead of strings for name parameters. Numerical Issue (RC.7), with 17 cases, involves failures in numerical computing, such as division by zero, overflow, and incorrect rounding. For example, Deepspeed issue #2090 treated microsecond timestamps as second units in flops profiling, causing metric inflation. Incorrect Assignment (RC.8), with 15 cases, includes wrong variable/value assignments. In Llama.cpp issue #2050, a tensor dimension index was misassigned during OpenCL kernel configuration.
Incorrect Synchronization (RC.9), with 3 cases, is crucial to the efficiency of LLM inference engines, indicating parallel execution flaws. For instance, vLLM issue #6449 failed to synchronize ‘RNG’ seed states across pipeline-parallel workers.
{mdframed}
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Finding 10: The complex functionality of LLM inference engines contributes to 41% of bug root causes. The diversity of nine failure modes highlights the need for formal specification checking, comprehensive unit testing, and runtime validation across components of engines.
5.1.4. [RD] Environment
Engines should adapt to various environments, which is a major source of bugs, with 301 cases (29% of total), covering three leaf categories.
Incompatible Backend (RD.1), with 138 cases, indicates errors in implementations for specific backends. For instance, in Llama.cpp issue #1737, the memory copy operation of ‘ggml_view_3d’ function encounters a crash due to the incompatibility of Metal backend. The vast number of RD.1 highlights that the compatibility of various backends and devices is the major challenge for LLM inference engines implementations. Incompatible Version (RD.2), with 98 cases, indicates incompatibilities arising from version changes in APIs. For example, in Mlc-llm issue #2212, an incorrect TVM version led to an API call error because the TVM API had been updated. This emphasizes the need for careful management of version and API changes in LLM inference engines. Dependent Module Issue (RD.3) involves errors or omissions in dependent libraries. For instance, in vLLM issue #8107, the docker image of the engine lacked the ‘timm’ dependency package. We observe 65 cases linked to this root cause, underscoring the challenge of managing extensive dependencies within the engine.
{mdframed}
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Finding 11: Environmental compatibility poses significant challenges for LLM inference engines, accounting for 29% cases, spanning three categories. This underscores the need for thorough compatibility testing, efficient version management, and reliable dependency management across various backends.
5.1.5. [RE] Resource
LLM inference engines are generally optimized for efficient resource management, where we observe 37 cases (4% of total) categorized into four leaf types.
Incorrect Resource Allocation (RE.1), with 14 cases, implies errors in allocating necessary resources. For example, in Llama.cpp issue #52, a miscalculation in memory allocation led to an out-of-memory error despite available free memory. Incorrect Cache Management (RE.2), with 11 cases, involves flawed implementations in cache management. For instance, issue #3825 in Llama.cpp demonstrates an incorrect shift operation within the KV cache. Incorrect Multi-device Management (RE.3), with 8 cases, deals with errors in handling multiple devices. In vLLM issue #7472, the engine failed to detect varying CUDA compute capabilities across GPUs, causing misallocation in multi-GPU setups. Incorrect Resource Deallocation (RE.4), with 4 cases, includes errors due to improper resource release. An example is TensorRT-llm issue #1190, where resources were potentially released multiple times in IPC environments. RE.4 requires particular caution because it could lead to severe consequences like memory corruption or security vulnerabilities.
{mdframed}
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Finding 12: Resource management bugs, though only 4%, are critical due to risks like memory leaks. We identify four leaf root causes, highlighting the need for carefully checking on resource allocation and release, cache management, and multi-device management.
5.2. Relationship Between Symptoms and Root Causes
Table 4. Relationship between bug symptoms and root causes.
Crash Unexpected Feature Abnormal System Silent
Output Failure Perform. Hang Error
Functionality 245 88 52 28 13 8
Environment 237 15 24 11 15 0
Configuration 153 26 33 14 14 4
In-/Output 16 13 9 1 2 0
Resource 22 3 3 7 1 1
To figure out the impact of each root cause and guide the location of root causes via symptoms, we explore the relationship between bug symptoms and root causes, as shown in Table 4.
Considering the consequence of root causes, Crash is the most common and harmful symptom across all five inner categories of root causes, making LLM services unavailable, highlighting the need for careful handling of all root causes. Additionally, secondary symptoms vary across root cause categories. Beyond crashes, bugs associated with Functionality and In-/output often lead to Incorrect Output. Environment and Configuration related issues are likely to cause Feature Failure due to incompatibilities. Resource-related issues can cause Abnormal Performance, because of the inefficient resource management. These findings align with expectations and indicate the correctness of our taxonomies.
{mdframed}
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Finding 13: All root cause categories are significant due to their consistently high likelihood of causing crashes.
For diagnosis of bug symptoms, first, the top three most common causes of Crash are bugs related to Functionality, Environment, and Configuration. This indicates that when a crash occurs, developers should first check functionality errors and test for reproducibility across environments and configurations to quickly identify the cause. Second, Feature Failure and System Hang are often equally linked to Environment, Functionality, and Configuration issues, so developers can consider these categories when observing the two symptoms. Third, for Incorrect Output and Silent Error, over half the bugs stem from Functionality errors, highlighting the need to check functionality implementation for these symptoms. Fourth, Abnormal performance is mainly due to Functionality and Configuration related issues.
{mdframed}
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Finding 14: We identify following diagnostic patterns: crashes, feature failures, and system hangs require checks on functionality, environment, and configuration; incorrect output and silent errors need functionality checks; and abnormal performance involves functionality and configuration checks.
(a) Symptoms.
(b) Root causes.
Figure 6. Correlation of characteristics across engines
6. RQ3: Commonality
To explore the commonality characteristics of bugs in different engines, following the previous work (Islam et al., 2019), we use the Pearson correlation coefficient ($r$) to analyze the correlation of bug symptoms and root causes among the engines. Our primary hypothesis is that the engines will be strongly correlated based on the distribution of bug symptoms and root causes as they are doing similar tasks.
Our analysis confirms that hypothesis as shown in Fig. 6. All the computed $r$ are greater than 0.5, indicating strong correlation (pea, 2025) in terms of the distributions of both symptoms and root causes. This results indicate that the distribution of bug symptoms and root causes is consistent across different engines, highlighting the generalizability of our findings. The correlation coefficients for bug symptoms are all above 0.9. This indicates that despite differences in targeted environments and implementations of LLM inference engines, bugs exhibit similar distributions and symptoms. This results further highlight that LLM app development requires a consistent bug tolerance method for LLM inference regardless of the engine used.
{mdframed}
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Finding 15: Bug symptoms and root causes across engines show a strong correlation ($r>0.5$), indicating common bug patterns despite implementation differences.
Correlation coefficients for root causes are slightly lower than those for bug symptoms. This is likely due to differences in engine implementations. Therefore, we further investigate the distribution of vulnerable components across these engines, as shown in Fig. 7, to explore potential implementation differences.
Figure 7. Vulnerable components across engines. Proportions exceeding 10% are annotated.
For engines designed for universal environments (including Llama.cpp and vLLM), the model conversion tool is the most bug-prone component. We observe a correlation coefficient of 0.64 between the vulnerable components of Llama.cpp and Mlc-llm, indicating their strong similarities. Specifically, in Llama.cpp, the most vulnerable components are model conversion tools and operators, accounting for 18% and 16% of bugs. Model conversion tools are bug-prone due to the need of Llama.cpp to convert various original model formats (such as Huggingface model format) into a unified format for compatibility with diverse backends and devices. Operators of Llama.cpp are also bug-prone because they are written manually from scratch in C++ for each backend (e.g., CPU, Vulkan, and Metal), increasing the risk of bugs. In Mlc-llm, model conversion tool issues account for 28% of bugs due to its reliance on TVM for compiling models to specific backends. This compilation process involves complex optimizations, making it particularly bug-prone.
For engines tailored to cloud, like vLLM and DeepSpeed, operators and model loaders are the most bug-prone components. The correlation coefficient between vLLM and DeepSpeed is 0.66, also indicating strong correlation in their vulnerable components. In vLLM, most bugs occur in the model loader and operators, accounting for 21% and 17%, respectively. While vLLM supports seamless compatibility with Hugging Face model formats, it is vulnerable during model loading (such as distributed loading across devices) and with operators optimized for cloud. Similarly, in DeepSpeed, bugs in operators and model loaders account for 27% and 16%, respectively.
TensorRT-LLM supports both cloud-based and edge server deployments. During the model loading phase, beyond model format conversion and quantization, TensorRT-LLM also requires specifying parallel and KV cache optimization strategies, which are typically handled during inference/serving in other engines. This complexity increases the bug rate for model conversion tools, accounting for 54% of all TensorRT-llm bugs.
{mdframed}
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Finding 16: The engines with the same targeted environment (e.g., universal or cloud) exhibit strong correlation ($r>0.5$) in bug-prone components.
7. Discussion
Implications. With the rising popularity of LLM inference engines, our study offers timely and practical insights for different stakeholders.
For inference engine vendors, our taxonomy in Finding 5.1 outlines 28 unique root causes, helping vendors understand and quickly address or avoid these issues. Findings 5.1.15.1.5 highlight areas that require special attention during implementations. For example, Finding 5.1.1 advises developers to concentrate on checking, processing, and error handling for character encoding, tokenizers, and user-specific formatting when dealing with inputs and outputs. While vendors might consider prioritizing different root causes, Finding 4 advises treating all bugs with equal seriousness, since any could pose a high risk of causing crashes. Additionally, developers might wonder which engine components are more prone to errors. This question is answered by Finding 7. For example, for engines targeting cloud environments, special focus should be on the implementation of model loaders and operators.
For LLM app developers, Finding 4.1 highlights the symptoms to consider for fault tolerance during the development of LLM apps, including crashes, unexpected outputs, feature failures, abnormal performance, system hangs, and silent errors. Furthermore, Findings 4.1.14.1.3 recommend stage-specific fault tolerance strategies for LLM app developers. For example, Finding 4.1.1 reveals that 97% of bugs appear as crashes during the setup phase, highlighting the importance of implementing crash fault tolerance during LLM apps setup. Moreover, Finding 6 shows that symptom distribution is consistent across engines, indicating that fault tolerance considerations are the same regardless of the engine used in LLM apps. Additionally, Finding 5.1.4 identifies environmental compatibility as a major (29%) error source, urging LLM app developers to clearly specify target environments and select suitable engines.
For researchers, Finding 4 is the first to empirically identify the features of bugs under LLM undetermined output tokens, including mis-encoded characters, incoherent semantics, inconsistent outputs, abnormal token lengths, and repeated substrings, providing insights for researchers to design oracles. Additionally, Finding 4.2 indicates that researchers should not rely solely on crashes as the oracle for LLM engine testing, as over 35% of bugs exhibit other symptoms. Furthermore, Findings 5.1.15.1.5 offer a foundation for developing more efficient testing methods. For instance, Finding 5.1.2 recommends focusing on configuration combinations, model compatibility, precision, and concurrency to efficiently test bugs under these conditions.
Feedback from the Community. To validate our findings, we have performed a preliminary test on Llama.cpp using our findings. Specially, we employ Finding 5.1.2 to generate configurations composition automatically and Findings 4.2 and 4 to set the ‘crash’, ‘mis-encoded characters’, ‘incoherent semantics’, ‘inconsistent outputs’, ‘abnormal token lengths’, and ‘repeated substrings’ as test oracles. As a result, two bugs [^3] To maintain anonymity, the links of issues here will be added after the paper is accepted. have been confirmed and fixed by the core developers as of this writing. One bug is linked to unexpected output token length under the ‘-n’ configuration, and the other causes a crash under the ‘-ctv’ configuration. This result further demonstrates the validity of our findings.
Threats to Validity. First, our analysis focuses on five LLM inference engines, which may limit the generalizability of our findings to all LLM inference engines. To mitigate this threat, we carefully selected engines with high popularity, well-known providers, and diverse targeted environments. Second, our findings are based on reported issues from GitHub, potentially introducing bias. To mitigate this, we collected a large number of issues, specifically 929 bug-related issues, through careful manual inspection. Third, our methodology involves manual bug categorization, which may introduce bias. To mitigate this, three authors with at least over two years of LLM apps development participate in taxonomy construction, resolving inconsistencies through discussion until reaching an agreement.
8. Related Work
Empirical Study on Bugs in Machine Learning Libraries. Bugs in machine learning frameworks have long been a significant concern in software engineering (Guo et al., 2019; Chen et al., 2020; Zhang et al., 2018; Chen et al., 2023; Guan et al., 2023; Islam et al., 2019). Chen et al. (Chen et al., 2020) have conducted a comprehensive survey and analysis on Stack Overflow of developers’ questions about deploying machine learning models in the cloud, mobile, and browsers, establishing a taxonomy of specific deployment challenges. Chen et al. (Chen et al., 2023) have conducted an empirical study on four popular deep learning frameworks and measured the test coverage achieved by three testing techniques. While previous studies have investigated bugs in machine learning deployments across various platforms, they primarily focus on traditional machine learning models. Our work focuses on bugs in LLM inference engines, which have unique characteristics, such as undetermined output, high resource demands, and compatibility for diverse environments.
Testing of Machine Learning Libraries. Many studies have proposed different methods to test machine learning frameworks, including compilers (Liu et al., 2023, 2022; Ma et al., 2023; Luo et al., 2021) and libraries (Gu et al., 2022; Luo et al., 2021; Wang et al., 2020; Li et al., 2023; Pham et al., 2019; Yang et al., 2023; Wei et al., 2022; Deng et al., 2023, 2024). Liu et al. (Liu et al., 2023) have proposed NNSmith, which leveraged symbolic constraint solving and gradient-based search to generate diverse yet valid DNN test models for fuzzing deep learning compilers. Additionally, some studies leverage large language models for testing deep learning libraries (Deng et al., 2023, 2024). TitanFuzz (Deng et al., 2023) leverages Large Language Models (LLMs) for generating valid input programs to fuzz deep learning libraries. Although existing work has explored various components of DL frameworks, no studies have specifically addressed the bugs in LLM inference engines. Our work provides an empirical foundation for future research on LLM inference engines. Moreover, our open-source dataset can serve as a data source for enhancing the LLM-based testing methods.
9. Conclusion
We present the first study on bugs in LLM inference engines, addressing a critical gap in understanding the symptoms, root cause, and commonality of bugs. By examining five well-known engines, we identify six bug symptoms, 28 root causes, and strong correlation of these characteristics across engines, offering insights into efficient bug detection and location. Our primary findings include that (1) Five factors are useful for oracle setting under undetermined LLM outputs; (2) Diverse oracles beyond crashes are needed for LLM inference engine testing; (3) We reveal actionable diagnostic patterns for all identified bug symptoms. Our research provides an empirical foundation for analyzing and testing LLM inference engines, with practical implications for different stakeholders.